# Chapter 4 Discrete subgroups of semisimple Lie groups, beyond lattices

### 26 Feb 2024 v1

@@ 5.1.3 Properties
    In this case, $\tau$ induces an embedding $\partial_{\infty}\tau_{i}:G^{\prime}/P^{\prime}\hookrightarrow\mathrm{Gr}_{i}(\mathbb{K}^{d})$ (where $G^{\prime}/P^{\prime}$ is the visual boundary of the symmetric space of $G^{\prime}$, see Remark 4.17) and the boundary map of $\rho_{0}$ is the composition of the boundary map $\partial_{\infty}\Gamma_{0}\to G^{\prime}/P^{\prime}$ of $\sigma_{0}$ (see Theorem 4.23) with $\partial_{\infty}\tau_{i}$. Moreover, by Fact 4.30 there is in that case a neighbourhood of $\tau\circ\sigma_{0}$ in $\mathrm{Hom}(\Gamma_{0},\mathrm{PGL}(d,\mathbb{K}))$ consisting entirely of $P_{i}$-Anosov representations (hence with finite kernel and discrete image — see Lemma 4.27.(3)).
5.1.4 Examples in higher real rank
@@
  • Section 3.2: It follows from the work of Benoist [15] that the ping pong groups of Claim 4.9 are quasi-isometrically embedded (see Remark 4.35) in $\mathrm{PGL}(d,\mathbb{R})$. They are in fact $P_{1}$-Anosov: see [41, 96].
  • Section 3.2: When they are defined by $B_{1}^{\pm},\dots,B_{m}^{\pm}$ which have pairwise disjoint closures, the Schottky groups in $\mathrm{PGL}(2n,\mathbb{K})$ of Nori and Seade–Verjovsky are $P_{n}$-Anosov (see [82]) and the crooked Schottky groups in $\mathrm{Sp}(2n,\mathbb{R})\subset\mathrm{SL}(2n,\mathbb{R})$ are $P_{1}$-Anosov (see [35]).
  • Section 3.3: By Facts 4.30 and 4.31, the Barbot representations of closed surface groups into $\mathrm{SL}(d,\mathbb{R})$ are $P_{1}$-Anosov. The Hitchin representations into $\mathrm{PSL}(d,\mathbb{R})$ are $P_{i}$-Anosov for all $11\leq i\leq d-$ 1: this is Labourie’s original result from [107], where he introduced Anosov representations. The maximal representations of closed surface groups into $\mathrm{SO}(2,n)\subset\mathrm{SL}(n+2,\mathbb{R})$ are $P_{1}$-Anosov (see [36, 81]). We refer to Figures 9 and 10 for some illustrations of boundary maps.
@@
  • Section 3.4: All (known) higher Teichmüller spaces consist of Anosov representations (see [22, 36, 81, 107]).
@@
    Not all Anosov representations of closed surfacegroups belong to higher Teichmüller spaces. For instance, the Barbot representations of $\pi_{1}(S)$ into $\mathrm{SL}(d,\mathbb{R})$ from Section 3.3 are $P_{1}$-Anosov, but their connected component in $\mathrm{Hom}(\pi_{1}(S),\mathrm{SL}(d,\mathbb{R}))$ contains representations that are not injective and discrete.
  • Section 3.4: The two known families of higher-dimensional higher-rank Teichmüller spaces that we mentioned for Gromov hyperbolic groups $\Gamma_{0}=\pi_{1}(M)$ are $P_{1}$-Anosov: for holonomies of convex projective structures, see [16], and for $\mathbb{H}^{p,q}$-convex cocompact representations, see [10] (case $q=1$) and [52, 53] (general case).
5.1.5 Interlude: eigenvalues and singular values
@@
As in Section 4.4, for any finitely generated group $\Gamma_{0}$, we choose a finite generating subset of $\Gamma_{0}$ and denote by $\mathrm{Cay}(\Gamma_{0})$ the corresponding Cayley graph, with its metric $\mathtt{d}_{\mathrm{Cay}(\Gamma_{0})}$. We denote by $X=G/K$ the Riemannian symmetric space of $G$, with its metric $\mathtt{d}_{X}$, and fix a basepoint $x_{0}\in X$. We denote the translation length as in (2).
@@
We now reinterpret Remark 4.35 using Notation 4.34. Let $\|\cdot\|_{\mathrm{Euc}}$ be the standard Euclidean norm on $\mathbb{R}^{d}$. For $G=\mathrm{PGL}(d,\mathbb{K})$ with $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$, we can take $K=\mathrm{PO}(d)$ or $\mathrm{PU}(d)$ and $x_{0}=eK\in G/K=X$, so that for any $g\in G$ lifting to $\hat{g}\in\mathrm{GL}(d,\mathbb{K})$ with $|\det(\hat{g})|=1$,
@@ 5.1.6 Characterisations
As in Remarks 4.24, in condition (3) we cannot remove the assumption that $\Gamma_{0}$ be Gromov hyperbolic. See [102, § 4.4] for further discussion.
5.1.7 Limit sets
@@
Fact 4.39.
Any $g\in\mathrm{PGL}(d,\mathbb{K})$ can be written as $g=\kappa\exp(a)\kappa^{\prime}$ for some $\kappa,\kappa^{\prime}\in K$ and a unique $a\in\mathfrak{a}^{+}$; the entries of $a$ are $\mu_{1}(\hat{g}),\dots,\mu_{d}(\hat{g})$ (see Notation 4.34) where $\hat{g}\in\mathrm{GL}(d,\mathbb{K})$ is any lift of $g$ with $|\det(\hat{g})|=1$.
@@
Let $z_{0}^{+}:=\mathrm{span}(e_{1},\dots,e_{i})\in G/P_{i}$ and $z_{0}^{-}:=\mathrm{span}(e_{i+1},\dots,e_{d})\linebreak\in G/P_{d-i}$, where $(e_{1},\dots,e_{d})$ is the canonical basis of $\mathbb{K}^{d}$. By Fact 4.39, for any $m$ we can write $g_{m}=\kappa_{m}\exp(a_{m})\kappa^{\prime}_{m}$ where $\kappa_{m},\kappa^{\prime}_{m}\in K$ and $a_{m}\in\mathfrak{a}^{+}$ is diagonal; the entries of $a_{m}$ are $\mu_{1}(\hat{g}_{m}),\dots,\mu_{d}(\hat{g}_{m})$ where $\hat{g}_{m}\in\mathrm{GL}(d,\mathbb{K})$ is any lift of $g_{m}$ with $|\det(\hat{g}_{m})|=1$.
@@
5.1.8 Cocompact domains of discontinuity
We end this section by briefly mentioning a generalisation to Anosov representations of a nice feature of rank-one convex cocompact representations. Namely, we have seen in Section 4.2 that for $\mathrm{rank}_{\mathbb{R}}(G)=1$, if $X=G/K$ denotes the Riemannian symmetric space of $G$, then any convex cocompact subgroup $\Gamma$ of $G$ acts properly discontinuously, with compact quotient, on the open subset $\Omega_{\Gamma}:=\partial_{\infty}X\smallsetminus\Lambda_{\Gamma}$ of $\partial_{\infty}X$. In that case, $\partial_{\infty}X$ is the unique flag variety $G/P$ of $G$ with $P$ a proper parabolic subgroup of $G$ (Remark 4.17).
@@ 5.1.8 Cocompact domains of discontinuity
When $b$ is skew-symmetric, i.e. $G=\mathrm{Sp}(2n,\mathbb{R})$, the set $\mathcal{L}$ is the space $\mathrm{Lag}(\mathbb{R}^{2n})$ of Lagrangians of $\mathbb{R}^{2n}$. In this setting, if $\rho(\Gamma_{0})$ is a “strong” crooked Schottky group as in Section 3.2, defined by $B_{1}^{\pm},\dots,B_{m}^{\pm}$ with pairwise disjoint closures, then the set $\Omega_{\rho(\Gamma_{0})}$ of Example 4.41 coincides with the set $\Omega=\mathrm{Int}(\bigcup_{\gamma\in\Gamma_{0}}\rho(\gamma)\cdot\mathcal{D})\subset\mathrm{Lag}(\mathbb{R}^{2n})$ of Section 3.2.
5.2 Anosov representations and convex cocompactness
Recall that when $\mathrm{rank}_{\mathbb{R}}(G)=1$, Anosov representations coincide with convex cocompact representations in the classical sense of Definition 4.12. When $\mathrm{rank}_{\mathbb{R}}(G)\geq 2$, Theorem 4.37 shows that Anosov representations have a number of similarities, in terms of their dynamics, with rank-one convex cocompact representations: see Remark 4.38. Another similarity, of a more geometric nature, is the existence of cocompact domains of discontinuity as in Example 4.41: given an Anosov representation $\rho:\Gamma_{0}\to G$, such a domain of discontinuity $\Omega\subset G/Q$ yields, by taking the quotient, a closed manifold $\rho(\Gamma_{0})\backslash\Omega$ locally modeled on $G/Q$, whose geometry can be quite interesting (see [132, § 5]). These manifolds $\rho(\Gamma_{0})\backslash\Omega$ do not satisfy any kind of convexity properties in general.
@@ 5.2 Anosov representations and convex cocompactness
5.2.1 Two attempts
Our starting point is the following special case of Fact 4.31.
Fact 4.42.
@@ 5.2.1 Two attempts
A natural first attempt would be to replace $\mathbb{H}^{n}$ by the Riemannian symmetric space of $G$. However, this turns out to be rather restrictive: Kleiner–Leeb [104] and Quint [119] proved that if $G$ is a real simple Lie group of real rank $\geq 2$, with Riemannian symmetric space $X=G/K$, then any Zariski-dense discrete subgroup of $G$, acting with compact quotient on some nonempty convex subset of $X$, is a cocompact lattice in $G$; in particular, $\Gamma$ is not Gromov hyperbolic and $\rho$ is not Anosov. Thus this approach does not provide a generalisation of Fact 4.42.
Instead, we make a second attempt by viewing $\mathbb{H}^{n}$ as a properly convex open set in projective space as in (1); we can then try to generalise Fact 4.42 by replacing $\mathbb{H}^{n}$ with any properly convex open subset $\Omega$ of $\mathbb{P}(\mathbb{R}^{n+1})$.
@@
Similarly to Fact 4.16, we can rephrase convex cocompactness in terms of some specific convex set in $\Omega$. Namely, define the orbital limit set $\Lambda^{\mathsf{orb}}_{\rho(\Gamma_{0})}(\Omega)$ of $\rho(\Gamma_{0})$ in $\Omega$ to be the set of accumulation points in $\partial\Omega$ of some $\rho(\Gamma_{0})$-orbit of $\Omega$; one easily checks that $\Lambda^{\mathsf{orb}}_{\rho(\Gamma_{0})}(\Omega)$ does not depend on the choice of $\rho(\Gamma_{0})$-orbit, because $\partial\Omega$ does not contain any nontrivial segments. Define the convex core $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}(\Omega)\subset\Omega$ of $\rho(\Gamma_{0})$ to be the convex hull of $\Lambda^{\mathsf{orb}}_{\rho(\Gamma_{0})}(\Omega)$ in $\Omega$ (i.e. the smallest closed convex subset of $\Omega$ whose closure in $\mathbb{P}(\mathbb{R}^{d})$ contains $\Lambda^{\mathsf{orb}}_{\rho(\Gamma_{0})}$). Similarly to Fact 4.16, for infinite $\Gamma_{0}$, the action of $\Gamma_{0}$ on $\Omega$ via $\rho$ is then convex cocompact if and only if it is properly discontinuous and $\rho(\Gamma_{0})\backslash\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}(\Omega)$ is compact and nonempty.
The following result is a generalisation of Fact 4.42 in this setting. It was first proved in [52] for representations $\rho$ with values in $\mathrm{PO}(p,q)$, and then in general in [53] and independently (in a slightly different form and under some irreducibility assumption) in [139]. See also [10, 16, 51, 112] for related earlier results.
@@
On the other hand, there exist $P_{1}$-Anosov representations that do not preserve any properly convex open subset of $\mathbb{P}(\mathbb{R}^{d})$: e.g. Hitchin representations (see Sections 3.3 and 5.1) into $\mathrm{PSL}(d,\mathbb{R})$ for even $d$.
@@
This actually yields a characterisation of $P$-Anosov representations into $G$ for any proper parabolic subgroup $P$ of any noncompact semisimple Lie group $G$, by considering an appropriate representation of $G$ to some large projective linear group. For instance, for $G=\mathrm{PGL}(d,\mathbb{R})$ and $P=P_{i}$ with $11\leq i\leq d-$ 1 as in Section 5.1, we can consider the natural representation $\tau_{i}:\mathrm{PGL}(d,\mathbb{R})\to\mathrm{PGL}(S^{2}(\Lambda^{i}\mathbb{R}^{d}))$ where $S^{2}(\Lambda^{i}\mathbb{R}^{d})$ is the second symmetric power of the $i$-th exterior power of the standard representation of $\mathrm{GL}(d,\mathbb{R})$ on $\mathbb{R}^{d}$. (For $i=1$, this identifies with $\tau:\mathrm{PGL}(d,\mathbb{R})\to\mathrm{PGL}(\mathrm{Sym}_{d}(\mathbb{R}))$ above.) Again, one can check that $\rho:\Gamma_{0}\to\mathrm{PGL}(d,\mathbb{R})$ is $P_{i}$-Anosov if and only if $\tau_{i}\circ\rho:\Gamma_{0}\to\mathrm{PGL}(S^{2}(\Lambda^{i}\mathbb{R}^{d}))$ is $P_{1}$-Anosov. Theorem 4.44 then implies the following.
@@
5.2.3 Sketch of proof of Theorem 4.44
@@
Suppose that $\Gamma_{0}$ is Gromov hyperbolic, that $\rho$ is $P_{1}$-Anosov with boundary maps $\xi_{1}:\partial_{\infty}\Gamma_{0}\to\mathbb{P}(\mathbb{R}^{d})$ and $\xi_{d-1}:\partial_{\infty}\Gamma_{0}\to\mathbb{P}((\mathbb{R}^{d})^{*})$, and that $\rho(\Gamma_{0})$ preserves a nonempty properly convex open subset $\Omega$ of $\mathbb{P}(\mathbb{R}^{d})$. If we have not chosen $\Omega$ carefully, then it is possible that $\partial\Omega$ contains segments or that the action of $\Gamma_{0}$ on $\Omega$ via $\rho$ is not convex cocompact. Therefore, we do not work with $\Omega$ itself, but consider instead the connected component $\Omega_{\max}$ of $\mathbb{P}(\mathbb{R}^{d})\smallsetminus\bigcup_{w\in\partial_{\infty}\Gamma_{0}}\xi_{d-1}(w)$ containing $\Omega$ (where we view each $\xi_{d-1}(w)$ as a projective hyperplane in $\mathbb{P}(\mathbb{R}^{d})$); it is $\rho(\Gamma_{0})$-invariant, open, and convex (not necessarily bounded) in some affine chart of $\mathbb{P}(\mathbb{R}^{d})$. Using Lemma 4.40, one can show that the action of $\Gamma_{0}$ on $\Omega_{\max}$ via $\rho$ is properly discontinuous, and that the set of accumulation points of any $\rho(\Gamma_{0})$-orbit of $\Omega_{\max}$ is $\xi_{1}(\partial_{\infty}\Gamma_{0})$.
Consider the convex hull $\mathcal{C}$ of $\xi_{1}(\partial_{\infty}\Gamma_{0})$ in $\Omega_{\max}$. One easily checks, using the transversality of $\xi_{1}$ and $\xi_{d-1}$, that $\xi_{1}(\partial_{\infty}\Gamma_{0})$ is not contained in a single supporting hyperplane to $\Omega_{\max}$ in $\mathbb{P}(\mathbb{R}^{d})$, and therefore that $\mathcal{C}$ is nonempty. Using the expansion property (6) of Theorem 4.37 for Anosov representations, a similar reasoning to the proof of (6$\Rightarrow$ (1) in Section 4.4 then shows that $\rho(\Gamma_{0})\backslash\mathcal{C}$ is compact: see [53, § 8].
By transversality of $\xi_{1}$ and $\xi_{d-1}$, there are no segments in $\partial\Omega_{\max}$ between points of $\xi_{1}(\partial_{\infty}\Gamma_{0})$. This makes it possible to “smooth out” $\Omega_{\max}$ to obtain a $\rho(\Gamma_{0})$-invariant properly convex open subset $\Omega^{\prime}\subset\Omega_{\max}$ containing $\mathcal{C}$ such that $\partial\Omega^{\prime}$ is $C^{1}$ and contains no segments: see [53, § 9]. The action of $\Gamma_{0}$ on $\Omega^{\prime}$ via $\rho$ is then convex cocompact. ∎

### 28 Mar 2024 v2

@@ 1 Anosov subgroups
    In this case, $\tau$ induces an embedding $\partial_{\infty}\tau_{i}:G^{\prime}/P^{\prime}\hookrightarrow\mathrm{Gr}_{i}(\mathbb{K}^{d})$ (where $G^{\prime}/P^{\prime}$ is the visual boundary of the symmetric space of $G^{\prime}$, see Remark 4.6) and the boundary map of $\rho_{0}$ is the composition of the boundary map $\partial_{\infty}\Gamma_{0}\to G^{\prime}/P^{\prime}$ of $\sigma_{0}$ (see Theorem 4.12) with $\partial_{\infty}\tau_{i}$. Moreover, by Fact 5.6 there is in that case a neighbourhoodof $\tau\circ\sigma_{0}$ in $\mathrm{Hom}(\Gamma_{0},\mathrm{PGL}(d,\mathbb{K}))$ consisting entirely of $P_{i}$-Anosov representations (hence with finite kernel and discrete image — seeLemma 5.3.(3)).
Examples in higher real rank
@@
  • Section 2: It follows from the work of Benoist [15] that the ping pong groups of Claim 3.5 are quasi-isometrically embedded (see Remark 5.11) in $\mathrm{PGL}(d,\mathbb{R})$. They are in fact $P_{1}$-Anosov: see [42, 98].
  • Section 2: When they are defined by $B_{1}^{\pm},\dots,B_{m}^{\pm}$ which have pairwise disjoint closures, the Schottky groups in $\mathrm{PGL}(2n,\mathbb{K})$ of Nori and Seade–Verjovsky are $P_{n}$-Anosov (see [84]) and the crooked Schottky groups in $\mathrm{Sp}(2n,\mathbb{R})\subset\mathrm{SL}(2n,\mathbb{R})$ are $P_{1}$-Anosov (see [36]).
  • Section 3: By Facts 5.6 and 5.7, the Barbot representations of closed surface groups into $\mathrm{SL}(d,\mathbb{R})$ are $P_{1}$-Anosov. The Hitchin representations into $\mathrm{PSL}(d,\mathbb{R})$ are $P_{i}$-Anosov for all $11\leq i\leq d-$ 1: this is Labourie’s original result from [109], where he introduced Anosov representations. The maximal representations of closed surface groups into $\mathrm{SO}(2,n)\subset\mathrm{SL}(n+2,\mathbb{R})$ are $P_{1}$-Anosov (see [37, 83]). We refer to Figures 9 and 10 for some illustrations of boundary maps.
@@
  • Section 4: All (known) higher Teichmüller spaces consist of Anosov representations (see [22, 37, 83, 109]).
@@
    Not all Anosov representations of closed surfacegroups belong to higher Teichmüller spaces. For instance, the Barbot representations of $\pi_{1}(S)$ into $\mathrm{SL}(d,\mathbb{R})$ from Section 3 are $P_{1}$-Anosov, but their connected component in $\mathrm{Hom}(\pi_{1}(S),\mathrm{SL}(d,\mathbb{R}))$ contains representations that are not injective and discrete.
  • Section 4: The two known families of higher-dimensional higher-rank Teichmüller spaces that we mentioned for Gromov hyperbolic groups $\Gamma_{0}=\pi_{1}(M)$ are $P_{1}$-Anosov: for holonomies of convex projective structures, see [16], and for $\mathbb{H}^{p,q}$-convex cocompact representations, see [10] (case $q=1$) and [53, 54] (general case).
Interlude: eigenvalues and singular values
@@
As in Section 4, for any finitely generated group $\Gamma_{0}$, we choose a finite generating subset of $\Gamma_{0}$ and denote by $\mathrm{Cay}(\Gamma_{0})$ the corresponding Cayley graph, with its metric $\mathtt{d}_{\mathrm{Cay}(\Gamma_{0})}$. We denote by $X=G/K$ the Riemannian symmetric space of $G$, with its metric $\mathtt{d}_{X}$, and fix a basepoint $x_{0}\in X$. We denote the translation length as in (2).
@@
We now reinterpret Remark 5.11 using Notation 5.10. Let $\|\cdot\|_{\mathrm{Euc}}$ be the standard Euclidean norm on $\mathbb{R}^{d}$. For $G=\mathrm{PGL}(d,\mathbb{K})$ with $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$, we can take $K=\mathrm{PO}(d)$ or $\mathrm{PU}(d)$ and $x_{0}=eK\in G/K=X$, so that for any $g\in G$ lifting to $\hat{g}\in\mathrm{GL}(d,\mathbb{K})$ with $|\det(\hat{g})|=1$,
@@
As in Remarks 4.13, in condition (3) we cannot remove the assumption that $\Gamma_{0}$ be Gromov hyperbolic. See [104, § 4.4] for further discussion.
Limit sets
@@
Fact 5.15.
Any $g\in\mathrm{PGL}(d,\mathbb{K})$ can be written as $g=\kappa\exp(a)\kappa^{\prime}$ for some $\kappa,\kappa^{\prime}\in K$ and a unique $a\in\mathfrak{a}^{+}$; the entries of $a$ are $\mu_{1}(\hat{g}),\dots,\mu_{d}(\hat{g})$ (see Notation 5.10) where $\hat{g}\in\mathrm{GL}(d,\mathbb{K})$ is any lift of $g$ with$|\det(\hat{g})|=1$.
@@
Let $z_{0}^{+}:=\mathrm{span}(e_{1},\dots,e_{i})\in G/P_{i}$ and $z_{0}^{-}:=\mathrm{span}(e_{i+1},\dots,e_{d})\linebreak\in G/P_{d-i}$, where $(e_{1},\dots,e_{d})$ is the canonical basis of $\mathbb{K}^{d}$. By Fact 5.15, for any $m$ we can write $g_{m}=\kappa_{m}\exp(a_{m})\kappa^{\prime}_{m}$ where $\kappa_{m},\kappa^{\prime}_{m}\in K$ and $a_{m}\in\mathfrak{a}^{+}$ is diagonal; the entries of $a_{m}$ are $\mu_{1}(\hat{g}_{m}),\dots,\mu_{d}(\hat{g}_{m})$ where $\hat{g}_{m}\in\mathrm{GL}(d,\mathbb{K})$ is any lift of $g_{m}$ with $|\det(\hat{g}_{m})|=1$.
@@
Cocompact domains of discontinuity
We end this section by briefly mentioning a generalisation to Anosov representations of a nice feature of rank-one convex cocompact representations. Namely, we have seen in Section 2 that for $\mathrm{rank}_{\mathbb{R}}(G)=1$, if $X=G/K$ denotes the Riemannian symmetric space of $G$, then any convex cocompact subgroup $\Gamma$ of $G$ acts properly discontinuously, with compact quotient, on the open subset $\Omega_{\Gamma}:=\partial_{\infty}X\smallsetminus\Lambda_{\Gamma}$ of $\partial_{\infty}X$. In that case, $\partial_{\infty}X$ is the unique flag variety $G/P$ of $G$ with $P$ a proper parabolic subgroup of $G$ (Remark 4.6).
@@
When $b$ is skew-symmetric, i.e. $G=\mathrm{Sp}(2n,\mathbb{R})$, the set $\mathcal{L}$ is the space $\mathrm{Lag}(\mathbb{R}^{2n})$ of Lagrangians of $\mathbb{R}^{2n}$. In this setting, if $\rho(\Gamma_{0})$ is a “strong” crooked Schottky group as in Section 2, defined by $B_{1}^{\pm},\dots,B_{m}^{\pm}$ with pairwise disjoint closures, then the set $\Omega_{\rho(\Gamma_{0})}$ of Example 5.17 coincides with the set $\Omega=\mathrm{Int}(\bigcup_{\gamma\in\Gamma_{0}}\rho(\gamma)\cdot\mathcal{D})\subset\mathrm{Lag}(\mathbb{R}^{2n})$ of Section 2.
2 Anosov representations and convex cocompactness
Recall that when $\mathrm{rank}_{\mathbb{R}}(G)=1$, Anosov representations coincide with convex cocompact representations in the classical sense of Definition 4.1. When $\mathrm{rank}_{\mathbb{R}}(G)\geq 2$, Theorem 5.13 shows that Anosov representations have a number of similarities, in terms of their dynamics, with rank-one convex cocompact representations: see Remark 5.14. Another similarity, of a more geometric nature, is the existence of cocompact domains of discontinuity as in Example 5.17: given an Anosov representation $\rho:\Gamma_{0}\to G$, such a domain of discontinuity $\Omega\subset G/Q$ yields, by taking the quotient, a closed manifold $\rho(\Gamma_{0})\backslash\Omega$ locally modeled on $G/Q$, whose geometry can be quite interesting (see [134, § 5]). These manifolds $\rho(\Gamma_{0})\backslash\Omega$ do not satisfy any kind of convexity properties in general.
@@ 2 Anosov representations and convex cocompactness
Two attempts
Our starting point is the following special case of Fact 5.7.
Fact 5.18.
@@
A natural first attempt would be to replace $\mathbb{H}^{n}$ by the Riemannian symmetric space of $G$. However, this turns out to be rather restrictive: Kleiner–Leeb [106] and Quint [121] proved that if $G$ is a real simple Lie group of real rank $\geq 2$, with Riemannian symmetric space $X=G/K$, then any Zariski-dense discrete subgroup of $G$, acting with compact quotient on some nonempty convex subset of $X$, is a cocompact lattice in $G$; in particular, $\Gamma$ is not Gromov hyperbolic and $\rho$ is not Anosov. Thus this approach does not provide a generalisation of Fact 5.18.
Instead, we make a second attempt by viewing $\mathbb{H}^{n}$ as a properly convex open set in projective space as in (1); we can then try to generalise Fact 5.18 by replacing $\mathbb{H}^{n}$ with any properly convex open subset $\Omega$ of $\mathbb{P}(\mathbb{R}^{n+1})$.
@@
Similarly to Fact 4.5, we can rephrase convex cocompactness in terms of some specific convex set in $\Omega$. Namely, define the orbital limit set $\Lambda^{\mathsf{orb}}_{\rho(\Gamma_{0})}(\Omega)$ of $\rho(\Gamma_{0})$ in $\Omega$ to be the set of accumulation points in $\partial\Omega$ of some $\rho(\Gamma_{0})$-orbit of $\Omega$; one easily checks that $\Lambda^{\mathsf{orb}}_{\rho(\Gamma_{0})}(\Omega)$ does not depend on the choice of $\rho(\Gamma_{0})$-orbit, because $\partial\Omega$ does not contain any nontrivial segments. Define the convex core $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}(\Omega)\subset\Omega$ of $\rho(\Gamma_{0})$ to be the convex hull of $\Lambda^{\mathsf{orb}}_{\rho(\Gamma_{0})}(\Omega)$ in $\Omega$ (i.e. the smallest closed convex subset of $\Omega$ whose closure in $\mathbb{P}(\mathbb{R}^{d})$ contains $\Lambda^{\mathsf{orb}}_{\rho(\Gamma_{0})}$). Similarly to Fact 4.5, for infinite $\Gamma_{0}$, the action of $\Gamma_{0}$ on $\Omega$ via $\rho$ is then convex cocompact if and only if it is properly discontinuous and $\rho(\Gamma_{0})\backslash\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}(\Omega)$ is compact and nonempty.
The following result is a generalisation of Fact 5.18 in this setting. It was first proved in [53] for representations $\rho$ with values in $\mathrm{PO}(p,q)$, and then in general in [54] and independently (in a slightly different form and under some irreducibility assumption) in [141]. See also [10, 16, 52, 114] for related earlier results.
@@
On the other hand, there exist $P_{1}$-Anosov representations that do not preserve any properly convex open subset of $\mathbb{P}(\mathbb{R}^{d})$: e.g. Hitchin representations (see Sections 3 and 1) into $\mathrm{PSL}(d,\mathbb{R})$ for even $d$.
@@
This actually yields a characterisation of $P$-Anosov representations into $G$ for any proper parabolic subgroup $P$ of any noncompact semisimple Lie group $G$, by considering an appropriate representation of $G$ to some large projective linear group. For instance, for $G=\mathrm{PGL}(d,\mathbb{R})$ and $P=P_{i}$ with $11\leq i\leq d-$ 1 as in Section 1, we can consider the natural representation $\tau_{i}:\mathrm{PGL}(d,\mathbb{R})\to\mathrm{PGL}(S^{2}(\Lambda^{i}\mathbb{R}^{d}))$ where $S^{2}(\Lambda^{i}\mathbb{R}^{d})$ is the second symmetric power of the $i$-th exterior power of the standard representation of $\mathrm{GL}(d,\mathbb{R})$ on $\mathbb{R}^{d}$. (For $i=1$, this identifies with $\tau:\mathrm{PGL}(d,\mathbb{R})\to\mathrm{PGL}(\mathrm{Sym}_{d}(\mathbb{R}))$ above.) Again, one can check that $\rho:\Gamma_{0}\to\mathrm{PGL}(d,\mathbb{R})$ is $P_{i}$-Anosov if and only if $\tau_{i}\circ\rho:\Gamma_{0}\to\mathrm{PGL}(S^{2}(\Lambda^{i}\mathbb{R}^{d}))$ is $P_{1}$-Anosov. Theorem 5.20 then implies the following.
@@
Sketch of proof of Theorem 5.20
@@
Suppose that $\Gamma_{0}$ is Gromov hyperbolic, that $\rho$ is $P_{1}$-Anosov with boundary maps $\xi_{1}:\partial_{\infty}\Gamma_{0}\to\mathrm{Gr}_{1}(\mathbb{R}^{d})=\mathbb{P}(\mathbb{R}^{d})$ and $\xi_{d-1}:\partial_{\infty}\Gamma_{0}\to\mathrm{Gr}_{d-1}(\mathbb{R}^{d})=\mathbb{P}((\mathbb{R}^{d})^{*})$, and that $\rho(\Gamma_{0})$ preserves a nonempty properly convex open subset $\Omega$ of $\mathbb{P}(\mathbb{R}^{d})$. Since $\Omega$ was chosen without care, it is possible that $\partial\Omega$ contains segments or that the action of $\Gamma_{0}$ on $\Omega$ via $\rho$ is not convex cocompact. Therefore, we do not work with $\Omega$ itself, but consider instead the connected component $\Omega_{\max}$ of $\mathbb{P}(\mathbb{R}^{d})\smallsetminus\bigcup_{w\in\partial_{\infty}\Gamma_{0}}\xi_{d-1}(w)$ containing $\Omega$ (where we view each $\xi_{d-1}(w)$ as a projective hyperplane in $\mathbb{P}(\mathbb{R}^{d})$); it is $\rho(\Gamma_{0})$-invariant, open, and convex (not necessarily bounded) in some affine chart of $\mathbb{P}(\mathbb{R}^{d})$. Using Lemma 5.16, one can show that the action of $\Gamma_{0}$ on $\Omega_{\max}$ via $\rho$ is properly discontinuous, and that the set of accumulation points of any $\rho(\Gamma_{0})$-orbit of $\Omega_{\max}$ is $\xi_{1}(\partial_{\infty}\Gamma_{0})$.
Consider the convex hull $\mathcal{C}$ of $\xi_{1}(\partial_{\infty}\Gamma_{0})$ in $\Omega_{\max}$. One easily checks, using the transversality of $\xi_{1}$ and $\xi_{d-1}$, that $\xi_{1}(\partial_{\infty}\Gamma_{0})$ is not contained in a single supporting hyperplane to $\Omega_{\max}$ in $\mathbb{P}(\mathbb{R}^{d})$, and therefore that $\mathcal{C}$ is nonempty. Using the expansion property (6) of Theorem 5.13 for Anosov representations, a similar reasoning to the proof of (6$\Rightarrow$ (1) in Section 4 then shows that $\rho(\Gamma_{0})\backslash\mathcal{C}$ is compact: see [54, § 8].
By transversality of $\xi_{1}$ and $\xi_{d-1}$, there are no nontrivial segments in $\partial\Omega_{\max}$ between points of $\xi_{1}(\partial_{\infty}\Gamma_{0})$. This makes it possible to “smooth out” $\Omega_{\max}$ to obtain a $\rho(\Gamma_{0})$-invariant properly convex open subset $\Omega^{\prime}\subset\Omega_{\max}$ containing $\mathcal{C}$ such that $\partial\Omega^{\prime}$ is $C^{1}$ and contains no segments: see [54, § 9]. The action of $\Gamma_{0}$ on $\Omega^{\prime}$ via $\rho$ is convex cocompact as desired. ∎